OpenStudy (jackfrost33):

Which of the following is a factor of x4 + 5x3 - 4x2 - 8x + 6

OpenStudy (jackfrost33):

how do you factor out that equation?

OpenStudy (tkhunny):

You LOOK for roots. If there are Rational roots for this thing they MUST be +/- 1, +/- 2, +/- 3, +/- 6. Those are your ONLY hope of factoring with Rational Numbers. Try them! Do you know the Rational Root Theorem?

OpenStudy (jackfrost33):

is it the same as remainder therem

OpenStudy (tkhunny):

No, you use the Remainder Theorem to determine if the possible rational roots ARE, in fact, roots. Calculate remainders until you get them all or you find a zero.

OpenStudy (jackfrost33):

i got 1

OpenStudy (tkhunny):

x = 1 is a rational root. This makes what a factor of the polynomial?

OpenStudy (jackfrost33):

(x+1)?

OpenStudy (tkhunny):

So close. x = 1 ==> x-1 = 0 (x-1)

OpenStudy (jackfrost33):

oh okay so the x=1 will equal x-1? so how would you get x+1

OpenStudy (tkhunny):

Why do you need x+1? It isn't a factor unless the original equation is written incorrectly. x = -1 ==> x+1 = 0